OpenStudy (anonymous):

reduced a problem to 8.66, teacher is looking for simplest radical form. Ideas?

jimthompson5910 (jim_thompson5910):

I can't say for sure, but that sounds like \[\Large \frac{\sqrt{3}}{2} \approx 0.866\] but you multiply by 10 to move the decimal over 1 spot \[\Large \frac{10\sqrt{3}}{2} \approx 8.66\] \[\Large 5\sqrt{3} \approx 8.66\] I would need to see the full problem to be sure though

OpenStudy (anonymous):

It is a triangle with a hypotenuse of 10 and shorter leg of 5, by Pythagorean theorem, the answer is 8.66. If you use 30-60-90 angle theorem the answer is 5√3. Sent these answers to the teacher and she asked for simplest radical form

jimthompson5910 (jim_thompson5910):

so \[\Large 5\sqrt{3}\] is the answer she wants

jimthompson5910 (jim_thompson5910):

if you use the pythagorean theorem you get a^2 + b^2 = c^2 x^2 + 5^2 = 10^2 x^2 + 25 = 100 x^2 = 100-25 x^2 = 75 x = sqrt(75) x = sqrt(25*3) x = sqrt(25)*sqrt(3) x = 5*sqrt(3) which is the same answer as stated above

OpenStudy (anonymous):

thank you!